22. Biphenyl, C12H10, is a nonvolatile, nonionizing solute that is soluble in benzene, C6H6. At 25
Answer #1
Since BP is non-volatile, it has negligible vapour pressure, so the only contribution to vapour pressure comes from benzene, which is volatile.

Raoult's Law states that the relative lowering of vapour pressure equals the mole fraction of solute in the solution.

BP has a RMM of 12*12+10*1=154 and benzene of 6*12+6*1=78, so 16.8g of BP contains 16.8/154=0.1090 mol and 33.7g of benzene contains 33.7/78=0.4320 mol.

The mole fraction of solute is 0.1090/(0.1090+0.4320 )=0.2014, so the VP is lowered by 0.2014*100.84=20.31 torr.

The VP of the solution is 100.84-20.31=80.53 torr.